### v2

# ConCodeEval: Evaluating Large Language Models for Code Constraints in Domain-Specific Languages

Abstract
Recent work shows Large Language Models (LLMs) struggle to understand natural language constraints for various text generation tasks in zero- and few-shot settings. While, in the code domain, there is wide usage of constraints in code format to maintain the integrity of code written in Domain-Specific Languages [^1] https://w.wiki/6jCH (DSLs) like JSON and YAML which are widely used for system-level programming tasks in enterprises. Given that LLMs are increasingly used for system-level code tasks, evaluating if they can comprehend these code constraints is crucial. However, no work has been done to evaluate their controllability over code constraints. Hence, we introduce ConCodeEval first-of-its-kind benchmark having two novel tasks for code constraints across five representations. Our findings suggest that language models struggle with code constraints. Code languages that perform excellently for normal code tasks do not perform well when the same languages represent fine-grained constraints.
ConCodeEval: Evaluating Large Language Models for Code Constraints in Domain-Specific Languages
**footnotetext: The first two authors contribute equally.
1 Introduction
Large language models (LLMs) have shown promising results (Brown et al., 2020) in generating coherent text and code in zero and few-shot settings, especially for resource-rich languages. However, their practical utility depends on their ability to follow instructions or constraints at various granularity, encompassing user and system requirements. Recent studies (Sun et al., 2023) indicate that LLMs struggle with fine-grained constraints in natural language prompts, especially in paraphrase generation and numerical planning tasks.
While previous work studied the controllability of language models (LMs) using fine-grained instructions in natural language (NL) format, the necessity to represent instructions in code format arises from a critical enterprise use case involving system-level programming. In enterprises, Listing 1: The JSON sample generated (highlighted in yellow) by the Granite 20B model does not adhere to the minContains and subsequent numerical constraints. Write a JSON sample with field values as per the JSON format schema given below. { "type": "array", "contains": { "type": "number", "multipleOf": 2.66, "exclusiveMinimum": 0.08231885995435284, "exclusiveMaximum": 5.1100233535478 }, "minContains": 7 } JSON sample: ‘‘‘ [2.66, 5.22, 8.88, 12.54, 16.2, 19.86, 23.52, 27.18] ‘‘‘ system-level code integrity is maintained through use-case-specific constraints, typically encoded using schemas. Schemas are instructions in structured code languages like JSON, YAML, XML, or Python to enforce constraints like data types and required fields, ensuring code integrity. For example, the schema in Listing LABEL:lst:task1 requires the sample to be an array of numbers. Each number must be a multiple of 2.66 and fall within the range defined by the exclusiveMinimum and exclusiveMaximum fields. Additionally, the array must contain at least seven elements. Following such schema constraints, developers write system-level code in Domain-Specific Languages (DSLs) in a format similar to JSON, YAML, or XML. DSLs are custom languages with specialized schemas and syntax suitable for a particular domain or application. These DSLs are common for tasks like data exchange and system configuration, such as in Kubernetes [^2] https://w.wiki/3kbZ. Writing DSL code requires deep domain expertise and a significant learning process for developers. This has led to a growing adoption of LLMs for system-level programming in several products such as Ansible Lightspeed [^3] https://developers.redhat.com/products/ansible/lightspeed.
Given the cruciality of factoring in schemas with LLMs, there is increasing interest in using constrained decoding for DSLs (Pimparkhede et al., 2024; Wang et al., 2024a). However, given its limitations (Appendix A.4), it is necessary to evaluate if LMs are cognizant of code constraints when directly presented as a part of the prompt. Therefore, we aim to study the controllability of LMs through two novel seed tasks: (i) Data as Code generation: valid sample generation factoring in constraints (ii) DSL validation: validate code against constraints. We evaluate two model families, Llama and Granite, ranging from 8B to 70B parameters, aligning with enterprise needs for system-level tasks, where open-source models provide an economical and transparent alternative to black-box models like GPT-4. Both tasks are highly motivated from research and enterprise use case point of view as detailed in Appendix A.5.
Our contributions are:
  1. 1.
    We introduce two novel NLP tasks for enterprise system-level code: code generation from fine-grained schema instructions and code validation against schemas. To the best of our knowledge, we are the first to evaluate LMs on these tasks.
  2. 2.
    A benchmark test set consisting of 602 schema samples, each containing multiple instructions. Each schema sample in our test set is represented in 5 different language formats (JSON, YAML, XML, Python, and NL).
  3. 3.
    Comparative and qualitative analysis of state-of-the-art code models for two novel tasks with different schema languages. Our findings show that language models best comprehend JSON schema (Tables 1, 2) and are agnostic of language proportions in pre-training data.
2 Data as Code Generation in DSL
Figure 1: Uniform trend of steep decline in performance across models for constraints positioned in the middle and beginning of the JSON schema context and output for task 1. We divide the schema in to $3$ equal portions Begin, Middle, and End, and put the violated constraints based on their locality into either of these three buckets.
Task description:
Given the schema, the generation task (see Listing LABEL:lst:task1) aims to produce a compliant data sample in DSL code format. We draw inspiration from several use cases (see Appendix A.5), including synthesizing schema-compliant data from LLMs’ parametric memory to train and evaluate smaller-sized models Song et al. (2020) and generating diverse sets of samples to be used in product test pipelines. For reliable DSL code generation, LLMs need to be schema-aware.
Dataset:
We synthetically prepare $602$ schemas for each of the $5$ representations having combinations of various constraints (Appendix A.6). First, we prepare JSON schemas using our combinatorial tool to generate a good mix of constraints. We then convert each JSON schema to XML and YAML schemas using automated tools. Further, we include resource-rich general-purpose language - Python using the Pydantic library generated using the Gemini-1.0-pro (Team et al., 2024) model as a code translation task. We extend our evaluation to NL representation generated using rule-based templates. We [^4] The schemas are manually validated by the paper’s authors. ensure equivalence of the generated schemas across languages. More details are in Appendix A.8.
Evaluation metric:
Each schema-compliant code output LLM generates is awarded one point where schema compliance is checked using a schema validator tool. We then utilize the accuracy metric (Gen Acc) over all samples to benchmark performance across the models. Additionally, we also report the percentage of samples generated with the invalid root data type (RTV%) and invalid samples (IS%) in Table 5. The root data type is the data type of the whole DSL sample. For example, the root data type of sample represented in Listing LABEL:lst:task1 is array. For IS and RTV metrics, the lesser the number, the better the performance.
Experimental setup:
We report greedy decoding results since it performed slightly better than beam search with a beam width of 3. More details including hyper-parameters in Appendix A.7.
Prompts:
We experiment with zero-shot and 3-shot prompting for each model. For 3-shot prompting, we identify errors from the zero-shot setting, then select shots similar to the most frequent errors. Examples of prompts are in Appendix LABEL:lst:task1. While we represent the zero-shot results in the main paper, few-shot results are in the Appendix (Table 2).
Results:
Among the $5$ schema representations studied, NL is best understood by models across all outputs. JSON and YAML schemas perform well for constraints in code despite their limited presence in pre-training data. Surprisingly, models struggle with constraints in Python, though it is the major portion of the pre-training data, and models also find XML schemas challenging. Notably, models did not exhibit a performance boost when schema and output representation languages were the same. As shown in Figure 1, models are sensitive to the locality of the constraints in the schema and struggle to factor in constraints present in the beginning and middle portions of the schema. Irrespective of their overall performance, models show a similar distribution of mistakes across all constraint types (see Table 4) and show improved performance in few-shot results (see Table 2) when shots relevant to the mistakes used. Among the two family of models, Llama3 70B performed the best followed by Granite 34B.
Output Representation
JSON YAML XML
Model Schema Gen Acc Val Acc Gen Acc Val Acc Gen Acc Val Acc
Llama3 8B JSON 28.2 56.0 29.2 45.0 7.9 47.0
Granite 8B 47.5 56.0 24.7 55.0 5.1 45.0
Granite 20B 50.4 52.0 37.7 44.0 10.1 53.0
Granite 34B 53.3 64.0 32.2 57.0 11.2 65.0
Codellama 34B 58.4 64.0 23.0 54.0 9.4 53.0
[…]
Llama3 8B XML 10.2 37.0 22.5 42.0 10.2 46.0
Granite 8B 18.9 47.0 12.1 44.0 8.4 52.0
Granite 20B 24.0 37.0 12.4 47.0 8.6 57.0
Granite 34B 18.7 68.0 18.1 58.0 8.6 58.0
Codellama 34B 8.8 46.0 14.2 46.0 8.6 50.0
[…]
Llama3 8B YAML 25.9 46.0 8.1 44.0 6.4 45.0
Granite 8B 47.0 47.0 15.7 50.0 8.6 44.0
Granite 20B 34.7 31.0 25.9 38.0 8.4 47.0
Granite 34B 52.1 68.0 26.4 61.0 8.6 58.0
Codellama 34B 48.0 59.0 27.9 53.0 9.1 58.0
[…]
Llama3 8B Python 13.7 43.0 10.2 42.0 11.6 43.0
Granite 8B 10.2 54.0 11.9 58.0 11.1 55.0
Granite 20B 14.6 45.0 11.7 67.0 7.3 44.0
Granite 34B 17.7 54.0 13.9 67.0 10.6 46.0
Codellama 34B 13.7 49.0 11.6 53.0 8.4 44.0
[…]
Llama3 8B NL 30.2 63.0 24.5 56.0 9.6 57.0
Granite 8B 52.3 59.0 42.1 61.0 11.1 58.0
Granite 20B 65.4 54.0 46.0 48.0 10.9 60.0
Granite 34B 69.7 55.0 55.1 46.0 10.9 56.0
Codellama 34B 60.4 57.0 40.6 57.0 8.69 50.0
[…]
3 DSL Validation
Task description:
There is a growing body of work Hada et al. (2024) on showing promising usage of LLMs as evaluators in many tasks. On similar lines, given the DSL sample and schema to validate, this task (see Listing LABEL:lst:task2) aims to determine the validity of the provided sample against the constraints through boolean question answering (QA). Also, the task is highly motivated from various use cases (see Appendix A.5) and throws light on LM’s understanding of the relation between requirements and output in various representations.
Dataset:
For each of the $602$ schemas across $5$ representations as described in Section 2, we generate $3$ data samples across JSON, XML, and YAML languages. First, these data samples are synthetically generated by parsing through the JSON schema by randomly pruning and selecting constraints, resulting in data samples of different lengths and constraints. We then convert the generated JSON data samples to equivalent YAML and XML formats. Dataset consists of $3076$ instances with $45\%$ of no and $55\%$ yes instances.
Listing 2: In the JSON sample, values for fields stingo and anisic do not adhere to schema constraints. But the Granite 34B model gives the incorrect answer (highlighted in yellow) as yes.
Question:
Does the JSON sample { "tamil": false, "baser": null, "anisic": 1906.34, "stingo": "officiis tellus. illum modi odit quas mattis nunc", "pigheadedness": 52.0 } adhere to all the constraints defined in JSON format schema
"type": "object",
"properties": {
"tamil": { "type": "boolean" },
"baser": { "type": "null" },
"anisic": { "type": "number", "multipleOf": 17.02 },
"stingo": { "type": "string", "maxLength": 20 },
"pigheadedness": {"type": "number", "exclusiveMinimum": 27.65410407394338, "maximum": 93.85523810367313 } },
"additionalProperties": false
Respond to yes or no.
Answer:
yes
Evaluation metric:
Since it is a boolean QA task, we use Macro average F1 (see Table 6) and Accuracy (Val Acc) as evaluation metrics (see Table 1).
Experimental setup:
The decoding strategy used here is similar to the generation task as mentioned in Section 2. More details in Appendix A.7.
Prompts:
The goal of this task is to answer with either yes or no. We experiment with zero- and few-shot prompting. With few shot prompting we provide one example each of yes and no answer. Results for few-shot prompting and examples of prompts are given in Appendix (Table 2).
Results:
Although NL representation excels in generation tasks, it degrades the validation performance of larger models like 70B. JSON, YAML and Python representations show effectiveness in one of the output formats however poorly perform for others. Like in task 1, models perform sub-optimally when both schema and output representations are same. In lines with task 1, XML stands as a challenging language for models. The Llama3 70B model perform best in validation like in task 1, with other models hovering around 50% Val Acc, likely reflecting random choice given the binary nature of the task. Smaller models, particularly the Llama3-8B with natural language representation, show notable improvement, as its pre-training is a combination of NL and code.
4 Related Work
Generation:
There is extensive work Muennighoff et al. (2024); Cassano et al. (2022) on evaluating capabilities of LLMs for various code tasks such as code completion, translation, etc for resource-rich languages like Python. Despite there being work Cassano et al. (2022) on multi-lingual code, there is scant attention to low-resource languages such as DSLs, though having crucial importance. In parallel, using LLMs as evaluators for low-resource languages is gaining interest, however limited, mainly focusing on languages like XML and INI Lian et al. (2024).
Controllability of LLMs:
While LLMs can handle coarse-grained constraints like sentiment, they struggle with fine-grained constraints, such as ending a text with a specific word Sun et al. (2023). Code schemas often require such fine-grained control, and to our knowledge, we are the first to explore LLM controllability for constraints in code.
5 Conclusion
We introduce two novel tasks - Data as Code generation and DSL validation to test the controllability of LLMs when constraints are in code format, which is a crucial use case for system-level programming tasks in enterprises. We evaluate LLMs over $5$ schema representations, including YAML, JSON, Python, XML, and NL, and $3$ output representations, including YAML, JSON, and XML. We conclude that model performance does not directly correlate with language’s portion in pre-training data. Models for task 1 best understand NL. However, this is not the case with the validation task. Interestingly, models underperformed when the schema and output representations were the same, and the locality of the constraints in the schema impacted their performance. Task 2 results show that most of the models, irrespective of their size, found it very challenging since they performed just above or under the 50% accuracy. We hope our work can help innovation in improving the capabilities of LLMs for such challenging use cases and serve as a valuable reference.
6 Limitations
While we explore the DSL validation task by generating yes or no, exploring the model’s reasoning can give a more comprehensive analysis of LLM’s understanding. Further, one can include more complex constraints in the future for general-purpose programming languages, like coding style constraints to write code along with natural language prompts and schema.
Ethics Statement
Custom-created datasets have been created synthetically using open-source tools. The language models, tools, and frameworks used for evaluation are open source and can be used without copyright issues.
Appendix A Appendix
A.1 Prompts
This section defines the prompts which are used for models. We report different prompts for every model tried here and report the best-performing prompt results. Generally, the model consists of a System Prompt followed by a prompt template specific to the model.
A.1.1 Common prompt
For zero shot inference, we use a common prompt as it is for all the models irrespective of the model’s prompt format and we observe best results for Task-1 with this prompt. The prompt is as follows.
Listing 3: common prompt
Write an {input_representation} sample with field values as per the {output_representation} format schema given below.
{schema}
{output_representation} sample:
A.1.2 Granite model family
The granite model generally follows the question-answering format. Task-1 prompts for granite family models are as follows.
System prompt:
System:
You are an intelligent AI programming assistant, utilizing a Granite code language model developed by IBM. Your primary function is to assist users in code explanation, code generation and other software engineering tasks. You MUST follow these guidelines: - Your responses must be factual. Do not assume the answer is yes when you do not know, and DO NOT SHARE FALSE INFORMATION. - You should give concise answers. You should follow the instruction and provide the answer in the specified format and DO NOT SHARE FALSE INFORMATION.
Prompt 2:
Listing 4: QA-prompt-1
{System prompt}
Question:
Write an {input_representation} sample with field values as per the {input_representation} format schema given below.
{schema}
Answer:
Prompt 3:
Listing 5: QA-prompt-2
{System prompt}
Question:
Write an {input_representation} sample with field values as per the {output_representation} format schema given below. Please wrap your code
answer using ‘‘‘
{schema}
Answer:
{output_representation} and {input_representation} are the variables where {input_representation} take the values JSON, YAML, XML, Python, and natural language. {output_representation} takes the values JSON, YAML, and XML.
A.1.3 Llama family
For codellama $34B$ model we wrap the common prompt in [INST] and [/INST] tags. For the llama3-8B model, we use the System prompt along with user tags  [^5] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct.
System prompt: You are a helpful, respectful, and honest assistant. Always answer as helpfully as possible, while being safe. Your answers should not include any harmful, unethical, racist, sexist, toxic, dangerous, or illegal content. Please ensure that your responses are socially unbiased and positive. If a question does not make any sense or is not factually coherent, explain why instead of answering something not correct. If you don’t know the answer to a question, please don’t share false information.
Other than this, similar to the granite family we try Question answering format and instruction to wrap the output in quotes (“‘).
Output Representation
JSON YAML XML
Model Schema Gen Acc Val Acc Gen Acc Val Acc Gen Acc Val Acc
Llama3 8B JSON 48.3 71.2 46.6 68.1 39.2 64.1
Granite 8B 51.2 69.2 52.3 66.1 47.8 65.8
Granite 20B 58.3 73.5 56.4 72.3 50.2 68.2
Granite 34B 66.3 76.2 64.5 75.4 51.3 73.2
Codellama 34B 65.1 75.1 63.4 73.2 50.6 71.2
[…]
Llama3 8B XML 46.6 65.8 42.3 63.4 36.6 60.1
Granite 8B 46.2 64.8 44.5 63.2 34.5 57.3
Granite 20B 50.4 66.7 48.2 64.1 36.4 56.1
Granite 34B 52.3 68.5 51.1 63.4 39.2 53.2
Codellama 34B 49.2 66.2 49.2 63.2 35.1 52.1
[…]
Llama3 8B YAML 46.7 67.2 45.3 64.2 43.5 63.2
Granite 8B 48.1 65.2 46.2 61.2 44.2 61.2
Granite 20B 52.3 68.9 49.7 66.7 47.8 65.1
Granite 34B 54.2 67.7 51.3 65.3 45.3 56.4
Codellama 34B 56.8 66.4 50.2 64.3 47.8 56.2
[…]
Llama3 8B Python 43.2 60.1 41.1 58.9 39.2 57.6
Granite 8B 45.1 60.5 46.7 59.4 37.4 56.0
Granite 20B 48.2 57.2 45.9 57.8 38.4 58.2
Granite 34B 50.6 59.2 47.1 55.6 41.3 57.3
Codellama 34B 47.2 56.4 45.3 57.2 39.2 55.1
[…]
A.2 Data statistics
This section represents schema length comparison for various languages.
Language Max schema length Average schema length
XML $3316$ $364.82$
JSON $1954$ $208.23$
YAML $1295$ $135.09$
Table 3: Schema length comparison
Constraint Llama3 8B Llama3 70B
type 302 49
exclusiveMinimum 18 44
multipleOf 170 42
minLength 47 21
contains 22 12
exclusiveMaximum 22 12
maximum 11 2
maxLength 7 19
additionalProperties 4 0
minimum 4 15
Table 4: Both the models, least and best performing irrespective of their performance show a similar distribution of mistakes for each constraint.
A.3 Few shot prompting results
Below are the results for Few-shot prompting. We experiment with 3 shot prompting. We observe that the majority of errors made by all the models are regarding short schema and the schema having root type of array as shown in sample LABEL:lst:task1. An example of a 3-shot prompt for a DSL generation task is shown below.
Few shot prompt
Listing 6: Few shot prompt
{System prompt}
Your task is to write a JSON sample with field values as per JSON format schema.
You are given a few examples demonstrating the same.
JSON format schema:
"type": "array",
"contains": {
"type": "boolean"
"minContains": 0
JSON sample:
[true, true, false]
JSON format schema:
"type": "string",
"format": "idn-email"
JSON sample:
"hchavezexample.org"
JSON format schema:
"type": "array",
"items": {
"type": "number",
"multipleOf": 5.82,
"exclusiveMinimum": 3.069158195370172
JSON sample:
A.4 Limitations of Constrained Decoding
This section outlines some common problems with constrained decoding and emphasises on why it cannot be a complete and a viable solution for factoring in schemas to generate compliant text using language models.
A.4.1 Inference Performance Bottleneck
Constrained decoding often negatively effects inference throughput widely mentioned as one of the major drawbacks in many works Wang et al. (2024b); Pimparkhede et al. (2024); Geng et al. (2023) due to involvement of token-level operations keeping track of the schema constraints and tokens generated so far. This latency can be a factor of the complexity of the schema, tokens generated so far, and the nature of the constrained decoding implementation. Further, advances such as batched inference  [^6] https://github.com/microsoft/batch-inference are not yet there for constrained decoding limiting their scalability and practical use.
A.4.2 Complex Engineering Effort
Implementing a constrained decoding system can involve instrumenting at the decoding phase of the language model while keeping track of the tokens generated so far and structured schema adherence which can involve implementation specific to a schema representation and may not be possible to generalize to any schema representation. For instance, most of the openly available constrained decoding systems  [^7] https://github.com/outlines-dev/outlines have limited support and not generalized to various schemas such as XML and output formats such as YAML and others. It is worthwhile to note that some approaches tend to convert scehmas to context free grammars, however, this approach is possible with common schema representations such as Python pydantic. Additionally, implementing such a system requires deep domain expertise.
A.4.3 Model Performance Bottleneck
LLMs have multiple failure modes that can likely be triggered through constrained decoding. Many works show that LLMs are sensitive to the text being fed into them and often deteriorate the model’s performance. Some examples being the reverse curse from Berglund et al. (2024), where LLM understanding "A is B" may not guarantee to learn "B is A". Another work Chen et al. (2024) shows that the order of the premises can have a substantial impact on the performance often affecting negatively. Such failures can be triggered when the natural flow of text generation is interrupted through constrained decoding over autoregressive generation. The problem can worsen when it involves mixed generation of structured output and unstructured NL text.
A.4.4 Limited Scope
Since constrained decoding needs access to the decoding phase of the language model, its often not possible to apply such decoding to hosted or gated LLM deployments.
Applying constrained decoding to some common use cases is not obvious. Given $n$ structured schemas from $s_{1}$ to $s_{n}$, unstructured NL text output as $k$ and structured output as $u$. Common use cases in natural language processing (NLP) such as summarization involve the following input-output relationship. For some arbitrary schema $i$, $s_{i}$rightarrow $u$. Further typical use cases involve factoring in $n$ multiple schemas and generate $m$ multiple structured outputs $(s_{1}...s_{n})$rightarrow $(k_{1}...k_{m})$.
Employing constrained decoding in such use cases is not viable since in the first use case, tasks that output $u$ cannot leverage constrained decoding and schema has to go into LLMs as input. When multiple schemas and structured outputs are involved, its not obvious to choose the right schema for decoding a particular structured output. Such common use cases substantially limit the scope of using constrained decoding.
Output Representation
JSON YAML XML
Model Schema IS (%) RTV (%) IS (%) RTV (%) IS (%) RTV (%)
Llama3 8B JSON 1.9 50.1 1.8 49.8 1.6 73.9
Granite 8B 2.9 31.0 2.8 57.3 17.1 70.26
Granite 20B 13.9 15.6 2.3 38.0 7.9 71.92
Granite 34B 2.6 23.5 2.6 48.6 4.1 73.08
Codellama 34B 3.6 17.9 1.8 51.4 3.7 71.12
Llama3 8B XML 12.9 64.1 6.1 52.8 4.8 73.5
Granite 8B 3.6 60.7 2.8 70.9 10.7 72.0
Granite 20B 2.1 53.3 1.9 73.9 12.2 70.5
Granite 34B 1.9 56.9 1.6 63.1 10.6 71.9
Codellama 34B 2.3 71.2 1.6 56.9 10.2 71.7
Llama3 8B YAML 1.3 53.3 3.1 62.4 0.4 74.5
Granite 8B 11.2 13.7 1.8 63.9 12.2 70.5
Granite 20B 1.6 39.8 1.4 56.6 10.7 72.0
Granite 34B 3.1 14.9 1.1 40.6 10.6 71.9
Codellama 34B 7.1 24.9 1.4 50.3 12.6 71.0
Llama3 8B Python 5.4 64.9 3.1 72.9 3.1 72.9
Granite 8B 2.4 73.0 2.3 70.9 10.7 72.71
Granite 20B 1.6 64.7 2.4 68.7 16.6 71.42
Granite 34B 2.6 61.2 2.4 66.9 8.9 69.35
Codellama 34B 5.6 65.1 2.9 64.1 14.1 69.1
Llama3 8B NL 5.8 50.4 3.4 54.1 5.6 73.9
Granite 8B 2.1 28.9 2.6 29.2 8.3 69.24
Granite 20B 2.9 0.6 2.8 30.2 7.97 69.24
Granite 34B 2.3 1.9 2.4 8.9 9.86 63.42
Codellama 34B 2.8 60.4 2.9 34.5 7.88 65.51
Table 5: Task 1 zero shot results having IS and RTV metric values. IS denotes the percentage of invalid samples and RTV denotes the percentage of sample root data type errors. For IS and RTV, the lesser the value better the performance.
A.5 Task Motivation
A.5.1 Data as Code Generation Task
This section describes use cases from enterprise and research point of view motivating data as code generation seed task in our study.
Enterprise Use Cases:
(i) Test case structured data generation to test application interfaces such as REST API endpoints. Often enterprises have large number of services exposing API endpoints that have to be tested and LLMs can be a drop in solution to generate test case data at scale. (ii) Structured configuration data generation for a particuilar use case and domain. Enterprise applications such as Kubernetes use DSLs for configuration and usage, preparing them require deep domain expertise and there is increasing motivation Pujar et al. (2023) to employ LLMs in enterprises to generate DSL code. (iii) Some more downstream tasks involving structured data such as forms and tables often represented in programmable format such as JSON can leverage LLMs to generate structured data to fill forms or tables leveraging the schema.
Research Use Cases:
(i) Since DSLs are typically low resource languages, LLMs are often employed Song et al. (2020) to synthesize data from LLMs to train and evaluate smaller-sized models. (ii) This task acts a as a seed for similar NLP use cases such as code translation.
        Output Representation
        JSON         YAML         XML
        Model         Schema         Macro-F1         Macro-F1         Macro-F1
        Llama3 8B         JSON         0.55         0.37         0.40
        Granite 8B         0.55         0.55         0.42
        Granite 20B         0.48         0.37         0.47
        Granite 34B         0.60         0.56         0.63
        Codellama 34B         0.64         0.53         0.50
        Llama3 8B         XML         0.44         0.35         0.41
        Granite 8B         0.45         0.44         0.50
        Granite 20B         0.24         0.45         0.56
        Granite 34B         0.52         0.47         0.39
        Codellama 34B         0.41         0.41         0.48
        Llama3 8B         YAML         0.38         0.40         0.40
        Granite 8B         0.45         0.50         0.44
        Granite 20B         0.24         0.31         0.45
        Granite 34B         0.52         0.55         0.47
        Codellama 34B         0.59         0.52         0.58
        Llama3 8B         Python         0.37         0.36         0.38
        Granite 8B         0.54         0.44         0.54
        Granite 20B         0.34         0.45         0.36
        Granite 34B         0.53         0.47         0.40
        Codellama 34B         0.48         0.45         0.46
        Llama3 8B         NL         0.63         0.55         0.57
        Granite 8B         0.45         0.51         0.39
        Granite 20B         0.53         0.45         0.57
        Granite 34B         0.45         0.46         0.38
        Codellama 34B         0.52         0.54         0.42
Table 6: Task 2 zero shot Macro-F1 scores. Task 2 is a binary classification task.
A.5.2 DSL Validation Task
This section describes use cases from enterprise and research point of view motivating DSL validation seed task in our study.
Enterprise Use Cases:
(i) Given the schema, employing LLMs to generate domain aware suggestions over the provided structured data which is not viable with traditional schema validators which only pinpoint syntactic errors and cannot provide semantic suggestions. Such as providing optimizations over the existing resource YAML in Kubernetes while complying with resource schema. (ii) In a assistive chat system, often the constraints are in NL representation from the user which are not machine readable and LLMs should be able to understand such constraints. (iii) Quick interoperability across different schema and data representation versions. Often in enterprises, schemas can be in a particular version not compatible with structured data’s version. For instance, the schema could be in an older JSON schema version such as Draft 0 and data in Draft 7, in such cases LLMs can come handy to perform validation at scale.
Research Use Case:
Understanding LLMs’ capability in validating the given structured data against the schema across representations can provide seed evidence for more complex tasks such as automatic fixing of data in compliance with given schema.
A.6 Schema Examples
This section provides schemas across $5$ representations from LABEL:lst:samplejson, LABEL:lst:sampleyaml, LABEL:lst:samplepython, LABEL:lst:samplexml and LABEL:lst:samplenl. All the schemas are equivalent in terms of constraints.
Listing 7: Sample schema using JSON Schema
"type": "object",
"properties": {
"footbaths": {
"type": "boolean"
"deluded": {
"type": "null"
"bravadoing": {
"type": "number",
"exclusiveMaximum": 5.131849487240756
"queintise": {},
"manucodia": {
"type": "number"
"antagonized": {},
"outbacker": {
"type": "number"
"sphenotripsy": {
"type": "boolean"
"hw": {
"type": "null"
"additionalProperties": true,
"required": []
Listing 8: Sample schema using YAML
additionalProperties: true
properties:
antagonized: {}
bravadoing:
exclusiveMaximum: 5.131849487240756
type: number
deluded:
type: null
footbaths:
type: boolean
hw:
type: null
manucodia:
type: number
outbacker:
type: number
queintise: {}
sphenotripsy:
type: boolean
required: []
type: object
Listing 9: Sample schema using Python
from pydantic import BaseModel, Field
class Schema(BaseModel):
footbaths: bool
deluded: None = Field(None, alias="null")
bravadoing: float = Field(..., exclusive_maximum=5.131849487240756)
queintise: None = {}
manucodia: float
antagonized: None = {}
outbacker: float
sphenotripsy: bool
hw: None = Field(None, alias="null")
Listing 10: Sample schema using XML
<?xml version="1.0" ?>
<all>
<type type="str">object</type>
<properties type="dict">
<footbaths type="dict">
<type type="str">boolean</type>
</footbaths>
<deluded type="dict">
<type type="str">null</type>
</deluded>
<bravadoing type="dict">
<type type="str">number</type>
<exclusiveMaximum type="float">5.131849487240756</exclusiveMaximum>
</bravadoing>
<queintise type="dict"/>
<manucodia type="dict">
<type type="str">number</type>
</manucodia>
<antagonized type="dict"/>
<outbacker type="dict">
<type type="str">number</type>
</outbacker>
<sphenotripsy type="dict">
<type type="str">boolean</type>
</sphenotripsy>
<hw type="dict">
<type type="str">null</type>
</hw>
</properties>
<additionalProperties type="bool">true</additionalProperties>
<required type="list"/>
</all>
Listing 11: Sample schema in NL
This is a JSON schema that defines the structure of an object. Heres a breakdown of the schema:
# **Top-level properties**
# * type: The type of the JSON data, which is an object (‘"object"‘).
# * properties: An object that defines the properties of the object.
# * additionalProperties: A boolean value that indicates whether additional properties not specified in the schema are allowed. In this case, it is set to True
* required: An empty array that specifies no properties are required in the object.
**Properties object**
The properties object defines the structure of each property in the object. Heres a brief description of each property:
footbaths: A boolean
deluded: A null
bravadoing: A number that must be strictly lesser than 5.131849487240756,
queintise: An object with no specific type or constraints.
manucodia: A number
antagonized: An object with no specific type or constraints.
outbacker: A number
sphenotripsy: A boolean
hw: A null
A.7 Hyperparameter details
We perform inference for all the models in $float16$ precision and a max new token limit of 1024 tokens. For beam search decoding, we use the beam width of 3.
A.8 Data Generation Scripts
A combinatorial data generation tool is created, which factors in constraints of interest, constraint-specific information, and combinatorial preferences, to generate the schemas. We also use openly available automatic lossless language to language translation tools for translation code from JSON to YAML etc. We plan to open-source all the scripts used for data preparation.